# Extreme time extrapolation capabilities and thermodynamic consistency of physics-inspired Neural Networks for the 3D microstructure evolution of materials

### v1

@@ IV Conclusions
This work shows that Convolutional Recurrent Neural Networks provide an excellent approach to approximate the microstructural evolution of materials undergoing the spinodal decomposition process in three dimensions. The use of fully convolutional and recurrent structures, together with a physics-inspired specialized layer, yields the accurate generation of spatiotemporal sequences at a fraction of the computational costs of the explicit method used to build the training set. Close correspondence for evolutions much longer than those provided in training has been demonstrated. Notably, the CRNN is found to properly predict the stationary states consistent with the learned equation, despite never being observed at train time.
@@
The data supporting the findings of this study are available from the authors upon reasonable request. Machine Learning code implementation is openly available at the following URL: https://github.com/dlanzo/CRANE.
@@ Appendix B Mesh rescaling effect on RMSE evaluation
We here report (Fig. 8) the RMSE plot of Fig. 5 as evaluated after downscaling the ground truth evolution on the same $64\times 64\times 64$ grid of collocation points used by the NN. Notice that the initial phases of the evolution now exhibit a lower RMSE value, due to the compensation of the resolution difference artifacts. In the subsequent phases of the evolution, however, values are only slightly affected, showing that the downscaling procedure is less relevant for coarser morphologies. As the RMSE on the $64\times 64\times 64$ grid is smaller than the one evaluated on the $128\times 128\times 128$ one, we report only the latter on the main text as it provides a stricter error measure.
@@
Appendix C $\bar{\delta}$ error measure

### v2

@@ IV Conclusions
This work shows that Convolutional Recurrent Neural Networks provide an excellent approach to approximate the microstructural evolution of materials undergoing the spinodal decomposition process in three dimensions. The use of fully convolutional and recurrent structures, together with a physics-inspired specialized layer, yields the accurate generation of spatiotemporal sequences at a fraction of the computational costs of the explicit method used to build the training set. Close correspondence for evolutions much longer than those provided in training has been demonstrated. Notably, the CRNN is found to properly predict the stationary states consistent with the learned equation, despite never being observed at train time. A main limitation of the present approach is that the trained model is only applicable to Cahn-Hilliard flow and new training, together with an ad-hoc training set, is required should the evolution law be changed. Additionally,should a non-conservative phenomenon be considered, e.g. dendritic growth or grain coarsening, a suitable modification of the physics-inspired layer should be implemented. Due to the flexibility of NN approaches and successful applications to other microstructure evolution problems Yang et al. (2021); Lanzoni et al. (2022); Peivaste et al. (2022); Fan et al. (2024), however, we expect that a suitably adapted version of the approach used in this work to be effective. On the other hand, if application to more challenging but related dynamical problems, such as evolution by surface diffusion, are to be tackled, the present model trained on Cahn-Hilliard flow could provide a good starting point for transfer learning procedures.
@@
The data supporting the findings of this study are available upon reasonable request from the authors. Machine Learning code implementation is openly available at the following URL: https://github.com/dlanzo/CRANE.
@@ Appendix B Mesh rescaling effect on RMSE evaluation
We here report (Fig. 8) the RMSE plot of Fig. 5 as evaluated after downscaling the ground truth evolution on the same $64\times 64\times 64$ grid of collocation points used by the NN. Notice that the initial phases of the evolution now exhibit a lower RMSE value, due to the compensation of the resolution difference artifacts. In the subsequent phases of the evolution, however, values are only slightly affected, showing that the downscaling procedure is less relevant for coarser morphologies and does not significantly impact the predictive capabilities of the NN model. As the RMSE on the $64\times 64\times 64$ grid is smaller than the one evaluated on the $128\times 128\times 128$ one, we report only the latter on the main text as it provides a stricter error measure.
@@
Appendix C Error measure
@@
Appendix S1 Performances assessment of NN performances if divergence layer is removed
We here report a comparison between the performances of the model considered in the main text and a variant that does not leverage the physics-inspired layer ("divergence" layer in Fig. 1(a) of the main text), showing that the first outperforms the latter one consistently, especially when very long sequences are considered. For the model version analyzed in this Section, the output layer is replaced by a logistic sigmoid activation function, following the approach used in 2D in Ref. Lanzoni et al. (2022). Consistently with the same work, we also add to the loss function a term acting as a (weak) penalty for the non-conservation of $\varphi$ to provide a fair comparison. All other hyperparameters and the training set are identical to the $\varphi$-conserving model. In the following, we will refer to the model considered in the main text as the "conservative model" and the variant not implementing the divergence layer described above as the "non-conservative model".
Figure S1: (a) Comparison between validation losses (calculated on the same set) for the conservative model discussed in the main text and a non-conservative one which does not implement the physics-inspired divergence layer. (b) Comparison between the finite difference ground truth, the conservative model prediction, and the non-conservative model one (evolution is the same as in Fig. 5 in the main text). (c) Comparison between the conservative and non-conservative model RMSE on the sequence of in panel (b).
Fig. S1(a) reports a comparison between the validation loss for the two models, calculated on the same set. A logarithmic scale is used to allow better inspection of small differences. Despite the oscillations, the conserving model consistently outperforms the non-conserving one.
Fig. S1(b) reports a comparison between the evolution obtained using the finite difference scheme and the prediction provided by both the conservative and non-conservative models. The initial condition is identical to the one of the evolution reported in Fig. 5(a) of the main text. Significant deviations are found in the sequence predicted by the non-conservative model in the late stages of the evolution, where a striking difference with respect to the expected layered configuration is found. It is also important to point out that error accumulation does not only affect the morphological prediction of the non-conservative model: while for short sequences in training only a fraction of percent in the deviation of the average value of $\varphi$ is observed, a high number of iterations such as the one reported in Fig. S1(b) leads to an $\approx 10\%$ deviation at $\tau=7.2\times 10^{4}$. This contrasts with the conservative model, in which the phase field is conserved up to numerical precision.
Finally, we also report in Fig. S1(c) the RMSE curve with respect to the ground truth along the whole temporal sequence. The equivalent error measure for the conservative model, already present in Fig. 5(b) of the main text is also reported for comparison. While in the first $\approx 200\tau$ the RMSE values are close, the non-conservative model error steeply increases when extrapolation with respect to the training set begins, saturating to a value almost twice the one observed for the conservative model. It can therefore be concluded that the use of the "divergence" layer in the CRNN structure is critical for accurate and physically consistent predictions in the long-time regimes.